BRICS Notes Series, Abstracts, 2003

December 17, 2003

This document is also available as PostScript and DVI.

References

[NS-03-3] PDF.

Luca Aceto, Zoltn sik, Willem Jan Fokkink, and Anna Inglfsdttir,

editors. Slide Reprints from the Workshop on Process Algebra: Open Problems

and Future Directions, PA ’03, (Bologna, Italy, 21–25 July, 2003), November 2003.

vi+138. Abstract: This volume contains reprints of the slides of most of

the talks that were delivered during the workshop on “Process Algebra: Open Problems and Future Directions” that was held in the period 21–25 July, 2003, at the University Residential Centre of Bertinoro, Forl, Italy.

[NS-03-2] PostScript, PDF, DVI.

Luca Aceto. Some of My Favourite Results in Classic Process Algebra. September 2003. 21 pp. To appear in the Bulletin of the EATCS, volume 81, October

2003. Abstract: This note collects some of the author’s favourite results

in classic process algebra, broadly construed, and points to some open problems in the field. It is hoped that this piece will offer a modest contribution to the healthy development of this area of research.

[NS-03-1] PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy,

Kurtz Alexander, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topology

in Concurrency Theory, GETCO ’03, (Marseille, France, September 6, 2003), August 2003.

vi+54. Abstract: This volume contains the preliminary proceedings of

GETCO ’03 held jointly with CMCIM 2003 in Marseille, France, on September 6, 2003, as a satellite workshop of CONCUR 2003.